# What Challenges Do Developers Face When Using Verification-Aware Programming Languages?

### v1

@@ I Introduction
  • An overview of practitioners first contact with VA languages.
@@ II-A Verification-Aware Languages Considered
Following a three-person analysis of the technologies listed, we arrived at the following set: Bandera [13], Dafny [34], Escher C Verifier [18], Eiffel [36], Frama-C [14], Java PathFinder [26], JML [33], mbeddr [51], Spec# [6], Spark Ada [5], VeriFast [28], VCC [12], Whiley [40], and Why3 [20].
@@ II-B Literature Review
After this process, nine of the 600 articles were considered relevant and were analyzed (see III-A).
@@ II-C Stack Overflow Q&A Study
This section details how we extract challenges from developer posts on Stack Overflow, a widely used Q&A platform where developers from diverse backgrounds openly discuss real-world programming problems at scale, making it a rich and authentic source for empirical studies [7, 3, 25, 1, 38, 2, 17, 41]. However, manually analyzing the vast amounts of data available in these sources may become infeasible, particularly as the dataset size increases. Therefore, we use topic modeling, a popular analytical tool for evaluating data, as an approach to identify groups of similar words, or topics, within the text [50]. One of the most popular algorithms for topic modeling is Latent Dirichlet Allocation (LDA) [8], an unsupervised machine learning algorithm that models automatically extracted topics in corpora. It is based on the idea that each document is a mixture of topics, and each topic is a mixture of words. LDA aims to discover how much of each topic is in each document and how much of each word belongs to a particular topic. The algorithm outputs probability distributions that quantify the proportion of each topic within a document and the association of each word with a given topic [47]. Several previous studies have used unsupervised machine learning, in particular LDA, for categorizing forum posts [7, 3, 25, 1, 38, 41, 2, 17].
Our analysis of forum posts using topic modeling follows from previous work and is divided into four main steps:
  1. 1.
    Data Collection: we collect data related to the considered VA languages from Stack Overflow.
  2. 2.
    Data Pre-processing: we implement various data pre-processing operations, such as removing all non-alphanumeric characters and markup language tags, including the ones that envelop code snippets and stop words, and standardization of words through Porter stemming [42] or lemmatization [19].
  3. 3.
    Model Training: LDA is trained using the dataset pre-processed in the previous step.
  4. 4.
    Tag Labeling: the tags are labelled, looking at words and topics belonging to each tag.
Next, we describe each step in detail.
@@
Three popular Stack Overflow post datasets are: SO official data dumps [21], SOTorrent [4], and SE Data Explorer [^2] https://data.stackexchange.com/). The first is distributed as a very large XML file, which we found difficult to process; the second is not updated since 2020; and the third is available through a human-centered interface, making it difficult to automate data collection. Therefore, we decided to use the official Stack Exchange API [^3] https://api.stackexchange.com/ as, similarly to SE Data Explorer, the API provides the newest data at the time of querying, but in a computer-friendly format.
We identified the tags we would use to query the API endpoint and settled with a variation of a list of VA languages presented in Section II-A. From our observations of the Stack Overflow tag list [^4] https://stackoverflow.com/tags, only hyphenated tags are supported, so “Escher C Verifier”, “Java PathFinder”, and “Spark Ada” had to be transformed to accommodate this requirement. In the final list mapping between official VA language names and the tag names, “Escher C Verifier”, “Java PathFinder”, and “Spark Ada” became “escher”, “pathfinder”, and “spark-ada”, respectively.
Moreover, we decided only to query the /questions endpoint because we are looking for challenges commonly found in questions but not in the follow-up interactions, i.e., answers and comments.
@@
  • __remove_code: Removes all tags and code snippets from the body of a given question, wrapping it into a div tag, and then performing a depth-first search into its children, removing all code and pre tags. Only the inner text of the other tags, such as anchor, strong, paragraph, etc, and text that is directly a child of a given tag is kept. The resulting text is then trimmed and joined with a white space. Based on [7], [3], [25], [38], [2], [17], and [41].
  • __gensim_preprocessing: Performs common pre-processing tasks using the functions strip_numeric, strip_multiple_whitespaces, and strip_punctuation from the Gensim [^5] https://radimrehurek.com/gensim/ library. This step removes digits from the initial string and multiple consecutive white spaces, leaving just one separating the words. Finally, it strips all the punctuation from the input string. Based on [3], [25], [38], [2], [17], and [41].
@@
  • __remove_stop_words: Filters out stop words using the NLTK stop words corpus (e.g., “a”, “an”, “the”, “of” and “and”). This is an approach used by [25], [1], and [41].
@@
  • __remove_handpicked: Removes words that, after examination, we considered as not adding value to the trained model. Based on [2] and [41].
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
Final Model. The selection of the final model configuration involved an evaluation across multiple training runs and criteria, which, due to space limitations, we briefly detail below. More details can be found in the supplementary material.
@@
For each combination of pre-processing pipelines and combination of hyperparameters, four models were trained. The performance of each combination (pipeline $\times$ combination of hyperparameters $\times$ 4 models) was assessed by calculating the average coherence and perplexity scores, as well as their standard deviations across the four models. This allowed to identify configurations that were not only high-performing but also consistent. Configurations with high standard deviations were considered unstable. Taking into consideration LDA’s non-deterministic nature, this strategy allows to detect if a score is exceptionally good or if the hyperparameters are able to make LDA consistently output good models.
@@
After performing open card sorting, we end up with the tag diversity shown in Figure 3. Despite being the most popular tag, the not-suited questions only comprise 19% of the overall questions against the 38% obtained in the previous section. Based on our analysis, we decided that this model was suited for naming and challenge identification.
@@ II-D1 Survey Design
We start the survey with a section about background information, where we ask about familiarity with VA languages and years of experience. We also ask about participants’ initial contact with VA languages and how they learned to program with them. We then attempt to validate the answers to RQ1. We use matrix-style questions with a 5-point Likert scale to ask questions about challenges related to Proof Elaboration and Environment Setup. We also ask participants about a subset of language-agnostic challenges identified in Section III-B and language-specific challenges with Dafny and Frama-C. We only showed these sections to participants who mentioned having experience with these languages.
We ask two open-ended questions about additional challenges not identified in the previous parts of the study and suggestions for improving the adoption of VA languages. Following these, we use a matrix-style closed-ended question to ask about the suggestions we developed based on the problems identified previously. Open-ended questions were presented before the closed-ended questions to prevent bias and allow the participants to think more freely without being conditioned by the options they already saw.
@@ II-D2 Recruitment
Our sample had a total of 31 participants. We recruited participants using our social media, professional network, and snowball sampling. We began by distributing the survey to 15 students of the 2023/2024 edition of [Formal Methods Course in First Authors’ Institution]. Using Proxycurl [^7] https://nubela.co/proxycurl/ [last access 29th July 2024], we gathered 28 profiles of LinkedIn users from the US with software and formal verification skills. We also contacted 16 industry professionals from our professional network and 26 Dafny contributors on GitHub, whose contact information was public. Additionally, we shared our survey in social media through X, formerly known as Twitter [^8] https://x.com/. We used snowball sampling by asking our contacts to distribute the survey to relevant members of their professional and student networks. We also distributed the survey through Frama-C’s mailing list. We divide participants by years of experience: more than three are considered experienced, all others are beginners.
@@
Some college, no degree 2 6.4%
@@
No, but have been in the past 2 6.4%
@@ II-D3 Quantitative Analysis
For our quantitative data, we use the Mann-Whitney U test to check whether the data from two groups, beginners and experienced, differ significantly. We chose a non-parametric test because our data does not follow a normal distribution. We use a Friedman test to compare answers to statements within the same group. If the Friedman test is significant ($p<0.05$), we then use the Wilcoxon signed-rank test to make pairwise comparisons of each statement and understand which particular statements are significantly different from the others. Finally, we analyze the frequency of each response to understand if the participants encountered the challenge. To analyze Likert scale data where we have both the Not Sure and Neutral options, we remove the Not Sure responses from the statistical test, assuming that these were due to the participant not understanding the statement.
@@ III-B Findings from Stack Overflow Q&A Study
The challenges in this class are related to the programming paradigm, so these are language agnostic. The first challenge is related to creating a proof strategy. In many question, authors could not even start the proof, not knowing what proof strategy to choose. This challenge was observed in nine questions.
@@
The challenges in this topic are language agnostic. The first class of challenges is related to dependency installation, which were identified in four questions. It includes challenges related to missing dependencies, incompatible dependency versions, and dependencies not being found. The other class of challenges is related to unsupported use cases. This challenge is represented by one question, in which the user wants to configure the version of function declarations used by the language server. The last class of challenges is missing installation instructions, also present in one question.
@@
The challenges in this topic are language-specific. We identified three challenges among Dafny programmers, all related to the modifies clause. The modifies clause specifies the set of memory locations that a method, iterator, or loop body may modify [35].
The first challenge is related to the user applying a modifies clause to a given memory location, holding a reference to an object within the scope of a method but only changing the properties of the given object inside a loop. Dafny assumes that the loop might change the object to which the memory location is pointing, thus making verification based on the previously set object properties impossible. The programmer must create an invariant saying that the memory location always points to the same object.
@@
Similarly to the previous topic, most of the challenges in this topic are related to features that are language-specific, related to Dafny and to the Frama-C WP Plugin, which uses ACSL as an annotation language.
@@
The following questions are considered language agnostic. In the first, the author wants to prove a lemma in Dafny. The original code of the author causes an internal error in Dafny 4.7.0. According to the accepted answer, an assertion was missing to make the verifier accept the proof. Because the need for the assertion is probably tied to the internal verification mechanism, we considered that the proof elicits a lack of language understanding, namely the underlying verification mechanisms. In the second question, the author is trying to understand the translation of mapping WhyML code into the SMT’s encoding to understand how the proofs for ACSL annotated programs are provided. This question is not a challenge but introduces a need for understanding the underlying verification mechanisms. The last question contains a simple syntax error where the author does not use the correct syntax to write comments in ACSL.
@@ III-C Findings from User Study
To assess differences between the seven conditions across all participants, we conducted a Friedman test. The results indicate a significant difference between the conditions, $chi^{2}(6)=27.79$, $p=0.0001$. To assess these differences, we conducted pairwise comparisons with a Wilcoxon signed-rank test with Bonferroni correction (see Table III). The results indicate that there were significant differences between $PE_{1}$ and $PE_{7}$ ($p=0.023$) and between $PE_{2}$ and $PE_{7}$ ($p=0.032$). No other comparisons reached statistical significance after the Bonferroni correction (all $p>0.05$). This means that $PE_{7}$, the challenge of expressing loop/recursion termination clauses is a significantly less important problem than $PE_{1}$ — designing proof strategies — and $PE_{2}$ — completing proof strategies.
@@
TABLE III: Pairwise Wilcoxon signed-rank test results with Bonferroni correction for multiple comparisons. *Significant at adjusted $p-value<0.05$.
@@
Here, we do not see a significant difference among experienced users. However, most developers, particularly beginners, struggled to understand how their high-level code is translated to the underlying SMT code, while more experienced developers felt confident in this area. When it came to translating proofs from other mediums (e.g., paper) to the verification language, most developers did not report significant difficulties. However, experienced developers showed relatively higher agreement with this challenge than beginners. Further investigation is needed to understand these trends fully.
@@
RQ1: VA language practitioners most commonly face challenges in designing and completing proof strategies. The only challenge where considerable differences between beginners and experienced users were identified is proving loop/recursion termination, though this issue was not common among most participants. On the other hand, this challenge is significantly less important than the challenges of designing and completing proof strategies. Language-specific semantics, such as frame specifications (modifies, loop assigns) and proof termination clauses, despite being identified in Stack Overflow, are not confirmed as challenges in our user study. While environment setup is less problematic, feature knowledge challenges were confirmed, with beginners struggling more with SMT translation, and experienced developers with translating proofs to the VA language. Fourteen new challenges were identified, with the most reported ones being “Tool Performance Issues”, “Poor Error Messages and Feedback”, “Complex Language Features and Low-Level Operations”, and “Specification and Annotation Challenges”.

### v2

@@ I Introduction
  • An overview of practitioners’ first contact with VA languages.
@@ II-A Verification-Aware Languages Considered
Following a three-person analysis of the technologies listed, we arrived at the following set: Bandera [14], Dafny [34], Escher C Verifier [19], Eiffel [36], Frama-C [15], Java PathFinder [26], JML [33], mbeddr [49], Spec# [7], Spark Ada [6], VeriFast [28], VCC [13], Whiley [39], and Why3 [20]. Although KeY [2] was considered a VA language, it was later removed as it introduced significant noise due to the name’s generality, including, e.g., results from security.
@@ II-B Literature Review
After this process, nine of the 600 articles were considered relevant and were analyzed (see III-A), highlighting a significant research gap and reinforcing the relevance of this study.
@@ II-C Stack Overflow Q&A Study
This section details how we extract challenges from developer posts on Stack Overflow, a widely used Q&A platform where developers from diverse backgrounds openly discuss real-world programming problems at scale, making it a rich and authentic source for empirical studies [8, 4, 25, 1, 3, 18, 40]. However, manually analyzing the vast amounts of data available in these sources may become infeasible, particularly as the dataset size increases. Therefore, we use topic modeling, a popular analytical tool for evaluating data, as an approach to identify groups of similar words, or topics, within the text [48]. One of the most popular algorithms for topic modeling is Latent Dirichlet Allocation (LDA) [9], an unsupervised machine learning algorithm that models automatically extracted topics in corpora. It is based on the idea that each document is a mixture of topics, and each topic is a mixture of words. LDA aims to discover how much of each topic is in each document and how much of each word belongs to a particular topic. The algorithm outputs probability distributions that quantify the proportion of each topic within a document and the association of each word with a given topic [45]. Several previous studies have used unsupervised machine learning, in particular LDA, for categorizing forum posts [8, 4, 25, 1, 40, 3, 18].
Our forum post analysis with topic modeling follows prior work and involves four steps: a) collecting Stack Overflow data on VA languages, b) pre-processing (cleaning, stop-word removal, stemming/lemmatization), c) training an LDA model, and d) labeling tags based on associated words and topics.
@@
Three popular Stack Overflow post datasets are: SO official data dumps [21], SOTorrent [5], and SE Data Explorer. [^2] https://data.stackexchange.com/ The first is distributed as a very large XML file, which we found difficult to process; the second is not updated since 2020; and the third is available through a human-centered interface, making it difficult to automate data collection. Therefore, we used the official Stack Exchange API [^3] https://api.stackexchange.com/ as, similarly to SE Data Explorer, it provides the newest data at the time of querying, but in a computer-friendly format.
We identified the tags we would use to query the API endpoint and settled with a variation of a list of VA languages presented in Section II-A. From our observations of the Stack Overflow tag list, [^4] https://stackoverflow.com/tags only hyphenated tags are supported, so “Escher C Verifier”, “Java PathFinder”, and “Spark Ada” had to be transformed to accommodate this requirement. In the final mapping between VA language names and the tag names, “Escher C Verifier”, “Java PathFinder”, and “Spark Ada” became “escher”, “pathfinder”, and “spark-ada”, respectively.
Moreover, we decided only to query the /questions endpoint because we are looking for challenges commonly found in questions but not in the answers and comments.
@@
  • __remove_code: Removes all tags and code snippets from the body of a given question, wrapping it into a div tag, and then performing a depth-first search into its children, removing all code and pre tags. Only the inner text of the other tags, such as anchor, strong, paragraph, etc, and text that is directly a child of a given tag is kept. The resulting text is then trimmed and joined with a white space. Based on [8], [4], [25], [3], [18], and [40].
  • __gensim_preprocessing: Performs common pre-processing tasks using the functions strip_numeric, strip_multiple_whitespaces, and strip_punctuation from the Gensim [^5] https://radimrehurek.com/gensim/ library. This step removes digits from the initial string and multiple consecutive white spaces, leaving just one separating the words. Finally, it strips all the punctuation from the input string. Based on [4], [25], [3], [18], and [40].
@@
  • __remove_stop_words: Filters out stop words using the NLTK stop words corpus (e.g., a, an, the, of, and). This is an approach used by [25], [1], and [40].
@@
  • __remove_handpicked: Removes words that, after examination, we considered as not adding value to the trained model  —  e.g., also, always, etc, like (full list available in the supplementary material). Based on [3] and [40].
@@
    baseline: This pipeline is the simplest and produces a list of individual words. The exact sequence of the steps is: __remove_code $\rightarrow\!$  __gensim_preprocessing  $\rightarrow\!$ __lemmatize_body $\rightarrow\!$ __make_lower_case  $\rightarrow\!$ __remove_stop_words.
@@
    baseline_bigram: Create bigrams from the list of words resulting from the baseline pipeline. Based on the work by [1]. The steps are: baseline $\rightarrow\!$__transform_bigrams.
@@
    rem_words_baseline: Similar to the regular baseline pipeline but incorporates the innovation of [3] and [40] of removing domain-specific words as well as an additional removal of words with less than three uses — since these words can not possibly form patterns — and less than two characters — since we considered that words with that size could not express meaningful value. The sequence is: baseline $\rightarrow\!$ __remove_handpicked $\rightarrow\!$__remove_uncommon $\rightarrow\!$ __remove_small.
@@
    rem_words_bigram: Based on the rem_words_baseline but, similarly to baseline_bigram, ends with a step that transforms the resulting list of words in a list of bigrams. Steps: rem_words_baseline $\rightarrow\!$__transform_bigrams.
@@
Final Model. We select the final model configuration through an evaluation across multiple training runs and criteria, which, due to space limitations, we briefly detail below. More details can be found in the supplementary material. [^6] https://figshare.com/s/bc220b6c300b46c97d89
@@
We trained a model for each combination of pre-processing pipelines and hyperparameters. The performance of each combination (pipeline $\times$ combination of hyperparameters $\times$ 4 models) was assessed by calculating the average coherence and perplexity scores, as well as their standard deviations across the four models. This allowed to identify configurations that were not only high-performing but also consistent. Configurations with high standard deviations were considered unstable. Taking into consideration LDA’s non-deterministic nature, this strategy allows us to detect if a score is exceptionally good or if the hyperparameters are able to make LDA consistently output good models.
@@
After performing open card sorting, we end up with the tag distribution shown in Figure 3. Despite being the most popular tag, the not-suited questions only comprise 19% of the overall questions against the 38% obtained in the initial training run. Based on our analysis, we decided that this model was suited for naming and challenge identification.
@@ II-D1 Survey Design
We start the survey with a section about background information, where we ask about familiarity with VA languages and years of experience. We also ask about participants’ initial contact with VA languages and how they learned to program with them. We then attempt to validate the answers to RQ1. We use matrix-style questions with a 5-point Likert scale to ask questions about challenges related to Proof Elaboration and Environment Setup. We also ask participants about a subset of language-agnostic challenges identified in Section III-B and language-specific challenges with Dafny and Frama-C. We only showed these sections to participants who mentioned having experience with these languages. This part of the survey mainly targets Dafny and Frama-C practitioners because over 70% of the questions found in Stack Overflow were related to these specific languages and some challenges are language dependent.
We ask two open-ended questions about additional challenges not identified in the previous parts of the study and suggestions for improving the adoption of VA languages. Then, we use a matrix-style closed-ended question to ask about the suggestions we developed a priori. We asked open-ended questions before the closed-ended to prevent bias and allow the participants to think more freely without being conditioned by the options they had already seen.
@@ II-D2 Recruitment
Our sample had a total of 31 participants. We recruited participants using our social media, professional network, and snowball sampling. We began by distributing the survey to 15 students of the 2023/2024 edition of the Formal Methods for Critical Systems course, an optional first-year Master’s level course from the Faculty of Engineering, University of Porto. Using Proxycurl, [^8] https://nubela.co/proxycurl/ we gathered 28 profiles of LinkedIn users from the US with software and formal verification skills, contacted 16 industry professionals from our professional network, and 26 Dafny contributors on GitHub, whose contact information was public. Additionally, we shared our survey on social media through X. [^9] https://x.com/ We used snowball sampling by asking our contacts to distribute the survey to relevant members of their professional and student networks. We also distributed the survey through Frama-C’s mailing list. We divide participants by years of experience: more than three are considered experienced, all others are beginners.
All participants were over 18 years old and provided informed consent after reviewing a form with the study’s purpose, procedures, and their rights. Participants were informed that they were not required to participate and could withdraw at any time without penalty. No personal or identifiable data were collected.
@@
Some college, no degree 2 6.5%
@@
No, but have been in the past 2 6.5%
@@ II-D3 Quantitative Analysis
For our quantitative data, we use the Mann-Whitney U test to check whether the data from two groups, beginners and experienced, differ significantly. We chose a non-parametric test because our data does not follow a normal distribution. We use a Friedman test to compare answers to statements within the same group. If the Friedman test is significant ($p<0.05$), we then use the Wilcoxon signed-rank test to make pairwise comparisons of each statement and understand which particular statements are significantly different from the others. Finally, we analyze the frequency of each response to understand if the participants encountered the challenge. To analyze Likert scale data, we remove the Not Sure responses from the statistical test, assuming that these were due to the participant not understanding the statement.
@@ III-B Findings from Stack Overflow Q&A Study
The challenges in this class are related to the programming paradigm, so these are language agnostic. The first challenge is related to creating a proof strategy. In many questions, authors could not even start the proof, not knowing what proof strategy to choose. This challenge was observed in nine questions.
@@
The challenges in this topic are language agnostic. The first class of challenges is related to dependency installation, which were identified in four questions. It includes challenges related to missing dependencies, incompatible dependency versions, and dependencies not being found. The other two classes of challenges are related to unsupported use cases and missing installation instructions. These challenges are represented by one question each.
@@
The challenges in this topic are language-specific, as we observed that languages where the user manually manipulates the memory, like C, have different challenges than garbage-collected ones, like Dafny.
We identified three challenges among Dafny programmers, all related to the modifies clause. The modifies clause specifies the set of memory locations that a method, iterator, or loop body may modify [35].
The first challenge is related to the user applying a modifies clause to a given memory location, holding a reference to an object within the scope of a method, but only changing the properties of the given object inside a loop. Dafny assumes that the loop might change the object to which the memory location is pointing, thus making verification based on the previously set object properties impossible. The programmer must create an invariant saying that the memory location always points to the same object.
@@
Similarly to the previous topic, most of the challenges in this topic are related to features that are language-specific, related to Dafny, and to the Frama-C WP Plugin, which uses ACSL as an annotation language.
@@
The following questions are considered language agnostic. In the first, the author wants to prove a lemma in Dafny. The original code of the author causes an internal error in Dafny 4.7.0. According to the accepted answer, an assertion was missing to make the verifier accept the proof. Because the need for the assertion is probably tied to the internal verification mechanism, we considered that the proof elicits a lack of language understanding, namely, the underlying verification mechanisms. In the second question, the author is trying to understand the translation of mapping WhyML code into the SMT encoding to understand how the proofs for ACSL annotated programs are provided. This question is not a challenge but introduces a need for understanding the underlying verification mechanisms. The last question contains a simple syntax error where the author does not use the correct syntax to write comments in ACSL.
@@ III-C Findings from User Study
To assess differences between the seven conditions across all participants, we conducted a Friedman test. The results indicate a significant difference between the conditions, $chi^{2}(6)=27.79$, $p=0.0001$. To assess these differences, we conducted pairwise comparisons with a Wilcoxon signed-rank test with Bonferroni correction (see Table III). The results indicate that there were significant differences between $PE_{1}$ and $PE_{7}$ ($p=0.023$) and between $PE_{2}$ and $PE_{7}$ ($p=0.032$). No other comparisons reached statistical significance after the Bonferroni correction (all $p>0.05$). This means that $PE_{7}$, the challenge of expressing loop/recursion termination clauses, is a significantly less important problem than $PE_{1}$ — designing proof strategies — and $PE_{2}$ — completing proof strategies.
@@
TABLE III: Pairwise Wilcoxon signed-rank test results with Bonferroni correction for multiple comparisons. *Significant at adjusted pvalue<0.05p-value<0.05.
@@
We do not see a significant difference among experienced users. However, most developers, particularly beginners, struggled to understand how their high-level code is translated to the underlying SMT code, while more experienced developers felt confident in this area. When it came to translating proofs from other mediums (e.g., paper) to the VA language, most developers did not report significant difficulties. Experienced developers showed relatively higher agreement with this challenge than beginners.
@@
RQ1: VA language practitioners most commonly face challenges in designing and completing proof strategies. The only challenge different between beginners and experienced users was proving loop/recursion termination, though this issue was uncommon and significantly less important than the challenges of designing and completing proof strategies. Language-specific semantics (e.g., frame specifications, termination clauses) appeared on Stack Overflow but were not confirmed as challenges in our user study. While environment setup is less problematic, feature knowledge challenges were confirmed, with beginners struggling more with SMT translation, and experienced developers with translating proofs to the VA language. We identified fourteen new challenges, the most reported ones being “Tool Performance Issues”, “Poor Error Messages and Feedback”, “Complex Language Features and Low-Level Operations”, and “Specification and Annotation Challenges”.